OpenStudy (anonymous):

In a study of 250 adults, the mean heart rate was 70 beats per minute. Assume the population of heart rates is known to be approximately normal with a standard deviation of 12 beats per minute. What is the 99% confidence interval for the mean beats per minute? 68.9 – 76.3 70 – 72 61.2 – 72.8 68 – 72

OpenStudy (anonymous):

Can anyone help? :(

OpenStudy (kropot72):

The 99% confidence interval is given by: \[\large \bar{x}-2.576\frac{\sigma}{\sqrt{n}},\ \bar{x}+2.576\frac{\sigma}{\sqrt{n}}\] Just plug in the given values.

OpenStudy (anonymous):

What goes where?

OpenStudy (anonymous):

Does 70 go into o and 12 go into n? or vice versa?

OpenStudy (kropot72):

If you are studying statistics you should be familiar with these symbols: \[\large \bar{x}=70\] \[\large \sigma=12\] \[\large n=250\]

OpenStudy (anonymous):

ok so